reinforced concrete design
Problem 2: Design of T-beam Calculate the reinforcing required for the T-beam system below 22 30 15 15 100 100 10 0 1 of 2
Comments
    Answer #1

    Given:

    Span=10 ft

    hf=8 in

    d=22 in

    L.L=400 lb/ft2

    Self weight of slab=(8 in/12x150) lb/ft3

    D.L of slab=100 psf

    effective flange width=span/4=60 in(assume it is least)

    Area of beam=8*60+22*15=810 in2

    D.L of beam=810x150/144=843.75 psf

    (a)

    Wu=1.2(D.L)+1.6(L.L)

    Wu=1.2(843.75)+1.6(400)

    Wu=1652.5 psf

    M=WL2/8

    M=1.6525x(20)2/8

    M=82.625 ft-kips

    (b)

    Effective flange width

    Span/4=20x12/4=60 in

    16hf+bw=16x8+15=143 in

    spacing=120 in

    The effective flange width is 60 in

    (c)

    Mn=M/(\phi)

    Mn=82.625/0.9=91.8 ft-kips

    Assume a lever arm of z=0.9d or z=d-hf/2

    z=0.90x22=19.8 or z=22-8/2=18 in

    AsFyZ=Mn

    As=91.8x12/(60x19.8)

    As=0.92 in2, which is less than the minimum area of steel

    Provide 2 no of #10 bars of area 2.53 in2

    equate the tension to compression

    0.85fcAc=AsFy

    0.85x4xAc=2.53x60

    Ac=44.64<240 in2

    a=44.64/60

    a=0.74<hf

    The minimum steel is 1.1 sq in providing 2.53 sq in of #10 no of 2 bars will be enough and a<hf

    (e)

    c=0.74/0.85

    c=0.87 in

    \epsilont=0.072>0.005

    The reduction factor is 0.9

    (f)

    Assume the cover of concrete to the T beam as 1.5 in. select a #4 diameter stirrup

    Check the spacing

    S=[bw-2(clear cover)-0.5(diameter of stirrup)-0.5(diameter of bar)]/2

    S=[15-2(1.5)-0.5(0.5)-0.5(1.25)]/2

    S=5.5 in

    Sketch:

    Sit 1、5 いSin

    (g)

    Moment capacity=0.90TZ

    =0.90AsFy(d-a/2)

    =0.90x2.53x60x(22-0.74/2)/12

    =246.25 ft-kips>91.1 ft kips, this is safe